A piece of paper area 60 cm square is divided in two pieces x and y so that the area of x is 2/3 of y. What is the area of x and y?
Found 2 solutions by rfer, solver91311:
Answer by rfer(16322)

x=2/3y
2/3y+y=60
y=60/(5/3)
y=60*3/5
y=36
2/3y=24

Answer by solver91311(24713)





But



So



Solve for and then calculate to get the value of

John

My calculator said it, I believe it, that settles it